a circle with radius 4 inches has a central angle of 45 degrees. What is the length of the inscribed arc?
Thanks

Found 2 solutions by Alan3354, fractalier:
Answer by Alan3354(69443)

a circle with radius 4 inches has a central angle of 45 degrees. What is the length of the inscribed arc?
-------------
Circumference = 2*pi*r
Arc = Circumference*(45/360)

Answer by fractalier(6550)

Length of an arc = [(theta)/360] * 2(pi)r = 45/360 * 8pi = pi inches